# Average Calibration Losses for Reliable Uncertainty in Medical Image Segmentation

### v2

@@ 1 Introduction
This work significantly extends our previous conference publication [3]. Specifically, we pair our marginal L1-ACE (mL1-ACE) auxiliary loss with a popular segmentation loss combination: Dice Similarity Coefficient (DSC) plus Cross-Entropy (CE), which generally yields better-calibrated predictions than DSC or CE alone [26]. We also extend mL1-ACE with a soft-binning approach, providing smoother gradients during training, and evaluate our method comprehensively on additional datasets.
@@ 2.3 Post-hoc Calibration Techniques
One approach to mitigating DNN miscalibration is to address it after training in a post-hoc fashion, without modifying network weights. A common approach is Temperature Scaling (TS) [14, 36], which multiplies logits by a scalar temperature to soften (temperature $>$ 1) or sharpen (temperature $<$ 1) predicted probabilities, and can be applied globally or per-class in segmentation tasks. Other methods include Platt scaling, isotonic regression, beta calibration, and Dirichlet calibration, with recent spatially aware extensions leveraging local neighbourhood information to further improve calibration [38]. These methods provide a means to improve calibration, without affecting the accuracy of the predictor, but very few parameters are leveraged for this calibration, limiting the ability to provide context-aware and instance-specific calibration.
@@ 2.4 Augmentation for Train-time Calibration
An alternative to post-hoc calibration is train-time calibration techniques, which leverages the high parameter count of DNN networks to achieve calibration. This can be performed indirectly, with techniques such as label smoothing, which softens hard one-hot labels, have been shown to reduce overconfidence and thereby improve calibration [39]. Additionally, data augmentation strategies, such as mixup [43, 40], also contribute to model calibration by forcing the network to produce softer, more uncertain outputs on interpolated samples. These methods modify the ground truth used in the optimisation process to mitigate overconfidence issues but they lack control over the actual calibration performance.
@@ 2.5 Auxiliary Calibration Losses in Classification
A direct approach to improve model calibration is to modify the training objective by incorporating an auxiliary loss that explicitly targets calibration. In image classification, several calibration-aware auxiliary losses have been proposed. For instance, Liang et al. [25] introduced a loss term based on the discrepancy between the predicted probabilities and the empirical accuracy (derived from the expected calibration error, ECE), penalising the model when reductions in the cross-entropy loss do not yield corresponding improvements in accuracy. According to Hebbalaguppe et al. [15], this DCA-based loss is purportedly limited by differentiability issues arising from discontinuities of the accuracy term due to the reliance on hard binning. To address this potential issue, Hebbalaguppe et al. [15] proposed the Multi-class DCA (MDCA) loss, which computes the calibration error by averaging over each class within a mini-batch rather than within fixed confidence bins. Moreover, Kumar et al. [23] introduced the Maximum Mean Calibration Error (MMCE) loss, a differentiable formulation that can be optimised jointly with the primary task. Other works have explored differentiable accuracy versus uncertainty calibration (AvUC) losses [22] and soft calibration objectives (S-AvUC) that employ continuous binning to produce differentiable estimates [19]. Recently, Bohdal et al. [7] developed a differentiable surrogate for expected calibration error (DECE) using soft binning and smooth approximations, although its reliance on meta-learning increases complexity and may limit wider applicability. As mentioned above, auxiliary losses have also been extended to canonical calibration [37]. Recent uncertainty-aware training strategies in medical image classification demonstrate that incorporating calibration-related loss terms can reduce miscalibration by a substantial margin while maintaining (or even improving) accuracy [10]. Dawood et al. showed that a Confidence Weight loss penalty on incorrect high-confidence predictions reduced the expected calibration error by 17–22% in cardiac MRI classification tasks, with slight accuracy gains.
@@ 2.6 Auxiliary Calibration Losses in Segmentation
In contrast to classification problems, segmentation tasks produce dense, per-pixel/voxel predictions, yielding a much larger number of confidence estimates per image. This can make calibration during mini-batch optimisation more effective, as the larger number of samples provides more stable and statistically meaningful estimates of calibration error for each forward pass. The use of auxiliary calibration losses in semantic segmentation is less common than in classification, yet it is equally critical in medical imaging applications. Standard segmentation losses, such as the Dice loss, have been shown to produce overconfident predictions [27], despite their effectiveness in handling class imbalance. To address this, Yeung et al. [42] proposed the DSC++ loss, which modifies the Dice loss by adjusting false positive and false negative penalties according to class frequency. Similarly, Neighbour-Aware Calibration (NACL) methods [30, 31] incorporate spatial consistency by penalising overconfident predictions in local neighbourhoods. Despite reported reduction in overconfidence with these approaches, there remains a need for a more controlled approach to ensure that model confidence estimates align more closely with observed accuracy while maintaining high segmentation performance.
@@ 4.1 Datasets
The ACDC dataset provides separate training and testing datasets, with $100$ cases in the training dataset and $50$ cases in the testing dataset. A training-validation split of $90$:$10$ was used, and the testing dataset was used as provided. For AMOS, only the CT images were used. The training dataset of 240 CT images was then split into training and validation with a ratio of $216$:$24$. The validation dataset was used as the testing dataset, as the testing dataset is not publicly available. For the BraTS dataset, the publicly available training dataset of $1251$ cases was split into training, validation and testing with a ratio of $1000$:$51$:$200$ cases, using the same testing split as our previous work [3]. For KiTS, the training dataset of $489$ cases was partitioned into training, validation and testing with a ratio of $352$:$39$:$98$.
@@ 4.2 Training and Evaluation
All training and evaluation was implemented via the MONAI framework [9]. Our open-source code can be found at https://github.com/cai4cai/Average-Calibration-Losses. A SegResNet model [32] was used, with 32 initial filters and 4 deep supervision layers, as employed by the KiTS 2023 competition winner [33]. We used the AdamW optimizer with a learning rate of $1\times 10^{-4}$, $2\times 10^{-4}$ for AMOS, and weight decay of $1\times 10^{-5}$. A warmup cosine scheduler was used, with a linear warmup period of 2 epochs with a warmup multiplier of 0.50, followed by cosine annealing for the remaining epochs. Training was run for 600 epochs for the ACDC and AMOS datasets and 1250 epochs for the BraTS and KiTS datasets. Validation was performed every 10 epochs, with the best model saved based on the validation Dice score. A fixed batch size of 2 was used, with the patch size maximised to fit within the GPU memory. All experiments were performed using an NVidia DGX cluster with V100 (32GB) and A100 (40GB) GPUs.
@@
For each dataset, three models were trained: (1) a baseline model using a Dice plus cross-entropy loss with a weight ratio of 1:1; (2) a model incorporating the baseline loss with an additional hard-binned auxiliary mL1-ACE loss (hL1-ACE), maintaining a ratio of 1:1:1 for Dice, cross-entropy, and ACE losses; and (3) a model incorporating the baseline loss with an additional soft-binned mL1-ACE loss (sL1-ACE), also maintaining a ratio of 1:1:1. For the auxiliary calibration losses, 20 bins were used to discretise the foreground probability space. For any classes with no ground truth present, the mL1-ACE loss was set to zero during optimisation.
The saved model with the highest validation Dice score was used for evaluation on the testing dataset. The Dice score, Average Calibration Error (ACE), Expected Calibration Error (ECE), and Maximum Calibration Error (MCE) were computed for each model, using both micro- and macro-averaging. For all calibration metrics, a fixed number of 20 bins was used, with hard-binning of predicted foreground probabilities. Additionally, reliability diagrams were generated for each testing case and a dataset reliability histogram for the whole dataset.
@@ 5 Results
We report results, averaged across classes, excluding background, and for BraTS and KiTS we report the average of the hierarchical evaluation classes (HEC). HECs reflect the labels used clinically and group together sub-regions into a larger class, such as whole tumour for BraTS, to create nested or composite labels. We evaluate the performance of the baseline model, trained with DSC and cross-entropy (CE), against models trained with mL1-ACE auxiliary losses, with both hard and soft binning. We compare the models across four datasets: ACDC, AMOS, BraTS, and KiTS. We achieve a Dice score of 0.880 on ACDC, 0.883 on AMOS, 0.905 on BraTS, and 0.859 on KiTS with the baseline model. This can be compared to state-of-the-art values of 0.927, 0.897, 0.915 and 0.887, respectively, as reported in the literature [17].
@@ 5.1 Tabular Results
We find that both hard- and soft-binned mL1-ACE outperform the baseline in terms of macro- and micro-averaged ACE and MCE across all datasets. In general, both variants are able to maintain DSC performance across datasets, though in some cases this may vary slightly. Notably, soft-binned mL1-ACE outperforms both hard-binned mL1-ACE and the baseline in terms of macro-averaged ACE (Table 1). Specifically, soft mL1-ACE improves macro-averaged ACE by 40% on ACDC, 19% on AMOS 30% on BraTS and 36% on KiTS compared to the baseline. For micro-averaged ACE, soft mL1-ACE provides improvements of 66% on ACDC, 52% on AMOS, and 58% on BraTS. However, on KiTS, hard-binned mL1-ACE achieves the largest improvement, with a 23% enhancement over the baseline, as shown in Table 1.
Table 1: Table showing the Dice Similarity Coefficient (DSC) and Average Calibration Error (ACE) metrics with macro- and micro-averaging for the baseline, hard mL1-ACE, and soft mL1-ACE loss function models across four testing datasets. Values show the average class metric, with BraTS and KiTS evaluated on the hierarchical class structure. Values in bold indicate the best performance for each dataset. Asterisks denote significance levels relative to the baseline: * p < 0.05, ** p < 0.01 (no symbol: non-significant). Standard deviation and statistical significance is not reported for micro-ACE as only one value is available for each dataset.
@@
Hard mL1-ACE ACDC 17 0.880 ± 0.035 0.102 ± 0.032∗∗ 0.069
@@ 5.4 Segmentation Probability Maps
We visualise the resultant probability maps from the models trained with the baseline and mL1-ACE losses in Fig. 5, for a given case and class from the testing dataset. These maps show the predicted foreground probability for each voxel for a given slice, with contours overlaid to show the relationship between the discrete prediction and the ground truth. We observe that the mL1-ACE losses result in a more diffuse probability distribution, with a softening of the class probabilities towards the decision boundary for the false positive (yellow) and particularly the false negative (red) regions of the segmentation. This indicates that the mL1-ACE losses are able to improve the calibration of the models, reducing the overconfidence in the model predictions.

### v3

@@ 1 Introduction
This work significantly extends our previous conference publication [3], introducing a soft-binned variant (sL1-ACE) with smoother gradients and substantially expanding our evaluation to four diverse medical imaging datasets. Additionally, we conduct sensitivity studies on bin count and loss weighting, include the use of micro-averaged metrics and update our network architecture. Throughout this work, we employ the popular Dice Similarity Coefficient (DSC) plus Cross-Entropy (CE) loss as our primary segmentation baseline [29]. Cross-entropy is a strictly proper scoring rule [16] and has been shown to yield better calibrated probabilities than DSC loss in medical image segmentation [30]. Consequently, adding a CE term to DSC (DSC+CE) is expected to improve calibration relative to pure DSC training.
@@ 2.1 Uncertainty Estimation in Medical Segmentation
It is important to note that methods modelling distributions over segmentations address a different aspect of uncertainty than confidence calibration. While distributional approaches capture epistemic and aleatoric uncertainty through multiple plausible outputs, confidence calibration focuses on ensuring that the predicted class probabilities from a single network output accurately reflect the likelihood of correctness. Both perspectives are valuable and complementary, however, our work addresses the latter, aiming to improve the reliability of softmax-derived confidence estimates.
@@ 2.2 Reliability Diagrams and Calibration Metrics
In the context of this work, we define calibration as confidence calibration: the alignment between the model’s predicted probability (confidence) and the true empirical accuracy. A model is considered well-calibrated if, for any given confidence level $p$, the proportion of correct predictions is approximately $p$.
Throughout this work, we use the terms confidence and predicted probability interchangeably, as both refer to the softmax output representing the model’s belief in a particular class. Similarly, accuracy and observed frequency are used interchangeably to denote the empirical correctness rate. This deliberate equivalence reflects the reformulation of the traditional confidence-accuracy framework into a probability-frequency perspective suited to dense prediction tasks.
@@ 2.3 Post-hoc Calibration Techniques
One approach to mitigating DNN miscalibration is to address it after training in a post-hoc fashion, without modifying network weights.
A common approach is Temperature Scaling (TS) [17, 39], which multiplies logits by a scalar temperature to soften (temperature >> 1) or sharpen (temperature << 1) predicted probabilities, and can be applied globally or per-class in segmentation tasks. Other methods include Platt scaling, isotonic regression, beta calibration, and Dirichlet calibration, with recent spatially aware extensions leveraging local neighbourhood information to further improve calibration [41, 13].
These methods provide a means to improve calibration, without affecting the accuracy of the predictor, but very few parameters are leveraged for this calibration, limiting the ability to provide context-aware and instance-specific calibration.
@@ 2.4 Augmentation for Train-time Calibration
An alternative to post-hoc calibration is train-time calibration techniques, which leverages the high parameter count of DNN networks to achieve calibration. This can be performed indirectly, with techniques such as label smoothing, which softens hard one-hot labels, have been shown to reduce overconfidence and thereby improve calibration [42]. Additionally, data augmentation strategies, such as mixup [47, 43], also contribute to model calibration by forcing the network to produce softer, more uncertain outputs on interpolated samples.
While these methods offer an implicit trade-off by reducing overconfidence but potentially degrading discriminative performance, they do not explicitly optimise a calibration metric and lack direct feedback on calibration quality during training.
@@ 2.5 Auxiliary Calibration Losses in Classification
A direct approach to improve model calibration is to modify the training objective by incorporating an auxiliary loss that explicitly targets calibration. In image classification, several calibration-aware auxiliary losses have been proposed. For instance, Liang et al. [28] introduced a loss term based on the discrepancy between the predicted probabilities and the empirical accuracy (derived from the expected calibration error, ECE), penalising the model when reductions in the cross-entropy loss do not yield corresponding improvements in accuracy. According to Hebbalaguppe et al. [18], this DCA-based loss is purportedly limited by differentiability issues arising from discontinuities of the accuracy term due to the reliance on hard binning. To address this potential issue, Hebbalaguppe et al. [18] proposed the Multi-class DCA (MDCA) loss, which computes the calibration error by averaging over each class within a mini-batch rather than within fixed confidence bins. Moreover, Kumar et al. [26] introduced the Maximum Mean Calibration Error (MMCE) loss, a differentiable formulation that can be optimised jointly with the primary task. Other works have explored differentiable accuracy versus uncertainty calibration (AvUC) losses [25] and soft calibration objectives (S-AvUC) that employ continuous binning to produce differentiable estimates [22]. Recently, Bohdal et al. [7] developed a differentiable surrogate for expected calibration error (DECE) using soft binning and smooth approximations, although its reliance on meta-learning increases complexity and may limit wider applicability. As mentioned above, auxiliary losses have also been extended to canonical calibration [40]. Recent uncertainty-aware training strategies in medical image classification demonstrate that incorporating calibration-related loss terms can reduce miscalibration by a substantial margin while maintaining (or even improving) accuracy [11]. Dawood et al. showed that a Confidence Weight loss penalty on incorrect high-confidence predictions reduced the expected calibration error by 17–22% in cardiac MRI classification tasks, with slight accuracy gains [11].
@@ 2.6 Auxiliary Calibration Losses in Segmentation
In contrast to classification problems, segmentation tasks produce dense, per-pixel/voxel predictions, yielding a much larger number of confidence estimates per image. This can make calibration during mini-batch optimisation more effective, as the larger number of samples provides more stable and statistically meaningful estimates of calibration error for each forward pass. The use of auxiliary calibration losses in semantic segmentation is less common than in classification, yet it is equally critical in medical imaging applications. Standard segmentation losses, such as DSC loss, have been shown to produce overconfident predictions [30], despite their effectiveness in handling class imbalance. To address this, Yeung et al. [46] proposed the DSC++ loss, which modifies the DSC loss by adjusting false positive and false negative penalties according to class frequency. Similarly, Neighbour-Aware Calibration (NACL) methods [33, 34] incorporate spatial consistency by penalising overconfident predictions in local neighbourhoods. Despite reported reduction in overconfidence with these approaches, there remains a need for a more controlled approach to ensure that model confidence estimates align more closely with observed accuracy while maintaining high segmentation performance.
@@ 4.1 Datasets
The ACDC dataset provides separate training and testing datasets, with $100$ cases in the training dataset and $50$ cases in the testing dataset. A training-validation split of 9090:$10$ was used, and the testing dataset was used as provided. For AMOS, only the CT images were used. The training dataset of 240 CT images was then split into training and validation with a ratio of 216216:$24$. The validation dataset was used as the testing dataset, as the testing dataset is not publicly available. For the BraTS dataset, the publicly available training dataset of $1251$ cases was split into training, validation and testing with a ratio of 10001000:5151:$200$ cases, using the same testing split as our previous work [3]. For KiTS, the training dataset of $489$ cases was partitioned into training, validation and testing with a ratio of 352352:3939:$98$.
@@ 4.2 Training and Evaluation
All training and evaluation was implemented via the MONAI framework [9]. Our open-source code can be found at https://github.com/cai4cai/Average-Calibration-Losses. A SegResNet model [35] was used, with 32 initial filters and 4 deep supervision layers, as employed by the KiTS 2023 competition winner [36]. We used the AdamW optimizer with a learning rate of $1\times 10^{-4}$, $2\times 10^{-4}$ for AMOS, and weight decay of $1\times 10^{-5}$. A warmup cosine scheduler was used, with a linear warmup period of 2 epochs with a warmup multiplier of 0.50, followed by cosine annealing for the remaining epochs. Training was run for 600 epochs for the ACDC and AMOS datasets and 1250 epochs for the BraTS and KiTS datasets. Validation was performed every 10 epochs, with the best model saved based on the validation DSC score. A fixed batch size of 2 was used, with the patch size maximised to fit within the GPU memory.
@@
For each dataset, we evaluated five configurations: (1) a DSC+CE baseline with a 1:1 weight ratio between DSC and cross-entropy; (2) a CE-only baseline trained with cross-entropy alone; (3) DSC+CE with an additional hard-binned mL1-ACE term (hL1-ACE) using a 1:1:1 weighting for DSC, CE, and ACE; (4) DSC+CE with a soft-binned mL1-ACE term (sL1-ACE), also with 1:1:1 weighting; and (5) a temperature-scaled DSC+CE model obtained via post-hoc temperature scaling.
Temperature scaling was applied post-hoc to the DSC+CE baseline by optimising a single temperature parameter on the validation set for one epoch using cross-entropy and the Adam optimiser (learning rate 0.001).
For the auxiliary calibration losses, 20 bins were used to discretise the foreground probability space. For any classes with no ground truth present, the mL1-ACE loss was set to zero during optimisation.
The saved model with the highest validation Dice Similarity Coefficient (DSC) was used for evaluation on the testing dataset. DSC, Average Calibration Error (ACE), Expected Calibration Error (ECE), and Maximum Calibration Error (MCE) were computed for each model, using both micro- and macro-averaging. For all calibration metrics, a fixed number of 20 bins was used, with hard-binning of predicted foreground probabilities. Additionally, reliability diagrams were generated for each testing case and a dataset reliability histogram for the whole dataset.
4.3 Sensitivity Studies
[…]
@@ 5 Results
We report results, averaged across classes, excluding background, and for BraTS and KiTS we report the average of the hierarchical evaluation classes (HEC). HECs reflect the labels used clinically and group together sub-regions into a larger class, such as whole tumour for BraTS, to create nested or composite labels. We evaluate the performance of the baseline model, trained with DSC and cross-entropy (CE), against models trained with mL1-ACE auxiliary losses, with both hard and soft binning. We compare the models across four datasets: ACDC, AMOS, BraTS, and KiTS. We achieve a DSC score of 0.871 on ACDC, 0.883 on AMOS, 0.905 on BraTS, and 0.859 on KiTS with the baseline model. This can be compared to state-of-the-art values of 0.927, 0.897, 0.915 and 0.887, respectively, as reported in the literature [20].
@@ 5.1 Tabular Results
We find that both hard- and soft-binned mL1-ACE outperform the baseline in terms of macro- and micro-averaged ACE and MCE across all datasets. In general, both variants are able to maintain DSC performance across datasets, though in some cases this may vary slightly. To contextualise these results, we additionally compare our proposed mL1-ACE losses against a CE-only baseline. Compared to DSC+CE, training with CE only yields modest improvements in macro-ACE (around 3–6% reduction across datasets) but consistently reduces DSC, illustrating the trade-off between segmentation accuracy and probabilistic calibration. Relative to CE only, Hard mL1-ACE further reduces macro-ACE by approximately 7–22% on three of the four datasets while maintaining DSC close to the DSC+CE baseline, and Soft mL1-ACE achieves even larger macro-ACE reductions of roughly 16–44% while preserving segmentation performance comparable to CE only (see Table 1). We also compare against post-hoc temperature scaling applied to the DSC+CE baseline. Temperature scaling improves macro-ACE by roughly 8–25% across datasets while maintaining DSC, confirming its effectiveness as a simple post-hoc calibration method. In comparison, Hard mL1-ACE attains calibration performance that is broadly comparable to temperature scaling with similar DSC, whereas Soft mL1-ACE further reduces macro-ACE by approximately 6–28% relative to temperature scaling, providing the strongest overall calibration improvements among the evaluated methods (Table 1).
Notably, soft-binned mL1-ACE outperforms both hard-binned mL1-ACE and the baseline in terms of macro-averaged ACE (Table 1). Specifically, soft mL1-ACE improves macro-averaged ACE by 46% on ACDC, 19% on AMOS, 30% on BraTS and 36% on KiTS compared to the baseline. For micro-averaged ACE, soft mL1-ACE provides improvements of 70% on ACDC, 52% on AMOS, and 58% on BraTS. However, on KiTS, hard-binned mL1-ACE achieves the largest improvement, with a 58% enhancement over the baseline, as shown in Table 1.
Table 1: Table showing the Dice Similarity Coefficient (DSC) and Average Calibration Error (ACE) metrics with macro- and micro-averaging for DSC+CE baseline, CE only, temperature scaling, hard mL1-ACE, and soft mL1-ACE loss function models across four testing datasets. Temperature scaling is applied post-hoc to the DSC+CE baseline. Values show the average class metric, with BraTS and KiTS evaluated on the hierarchical class structure. Values in bold indicate the best performance for each dataset. Asterisks denote significance levels relative to the DSC+CE baseline: * p < 0.05, ** p < 0.01 (no symbol: non-significant). Standard deviation and statistical significance is not reported for micro-ACE as only one value is available for each dataset.
@@
CE Only ACDC 17 0.868 ± 0.039 0.130 ± 0.028∗∗ 0.109
AMOS 22 0.872 ± 0.045∗∗ 0.103 ± 0.024∗∗ 0.053
BraTS 21 0.895 ± 0.130∗∗ 0.142 ± 0.063 0.047
KiTS 23 0.853 ± 0.139 0.160 ± 0.080 0.139
DSC + CE + Temp. Scaling ACDC 17 0.871 ± 0.037 0.102 ± 0.033∗∗ 0.073
AMOS 22 0.883 ± 0.044 0.093 ± 0.019∗∗ 0.033
BraTS 21 0.905 ± 0.107 0.135 ± 0.061∗∗ 0.033
KiTS 23 0.859 ± 0.144 0.145 ± 0.076∗∗ 0.088
DSC + CE + Hard mL1-ACE (ours) ACDC 17 0.870 ± 0.039 0.101 ± 0.029∗∗ 0.069
@@ 5.4 Segmentation Probability Maps
We visualise the resultant probability maps from the models trained with the baseline and mL1-ACE losses in Fig. 5, for a given case and class from the testing dataset. These maps show the predicted foreground probability for each voxel for a given slice, with contours overlaid to show the relationship between the discrete prediction and the ground truth. We observe that the mL1-ACE losses result in a more diffuse probability distribution, with a softening of the class probabilities towards the decision boundary for the false positive (yellow) and particularly the false negative (red) regions of the segmentation. This indicates that the mL1-ACE losses are able to improve the calibration of the models, reducing the overconfidence in the model predictions. All segmentation probability maps can be found on our public Github repository, including those from the 10 worst DSC cases for each dataset.
@@
5.5 Sensitivity Studies
Bin sensitivity. Figure 6 shows DSC and Macro-ACE when varying the number of bins from 5 to 100. Hard L1-ACE exhibits near-constant DSC and only minor variation in Macro-ACE across all bin counts. In contrast, Soft L1-ACE shows a reduction in DSC as the number of bins increases, while Macro-ACE remains relatively stable for both variants.
Loss weighting sensitivity. Figure 6 also reports the effect of varying the loss weighting $\lambda$ from 0.1 to 10.0. Increasing $\lambda$ reduces Macro-ACE for both Hard and Soft L1-ACE, indicating improved calibration. DSC for Hard L1-ACE remains stable over most of this range, with noticeable degradation only at the highest weighting. For Soft L1-ACE, DSC decreases more markedly as $\lambda$ increases, particularly for $\lambda\geq 1.0$.
Figure 6: Sensitivity study on the ACDC dataset for hard-binned (hL1-ACE) and soft-binned (sL1-ACE) mL1-ACE variants, showing segmentation performance (DSC) and calibration (Macro-ACE) as a function of the number of bins and the loss weighting $\lambda$.